### v1

# Detecting and Rectifying Noisy Labels: A Similarity-based Approach

Abstract
Label noise in datasets could damage the performance of neural net training. As the size of modern deep networks grows, there is a growing demand for automated tools for detecting such errors. In this paper, we propose post-hoc, model-agnostic error detection and rectification methods utilizing the penultimate feature from a neural network. Our idea is based on the observation that the similarity between the penultimate feature of a mislabeled data point and its true class data points is higher than that for data points from other classes, making the probability of label occurrence within a tight, similar cluster informative for detecting and rectifying errors. Extensive experiments show our method not only demonstrates high performance across various noises but also automatically rectifies these errors to improve the quality of datasets [^1] Our implementation is available at https://github.com/huutiendang/noise-detection-and-rectification.
1 Introduction
While the majority of knowledge in AI systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f. the LeCun’s cake). For Large Language Models (LLMs) such as GPTs Brown et al. (2020), LLaMA Touvron et al. (2023a, b), and Gemini Team et al. (2023), supervised learning accounts for only a small fraction of the total computation budget but has a significant impact on the models’ performance. Recent research (e.g. Zhou et al., 2023; Gunasekar et al., 2023) finds that high quality training data significantly improves performance while reducing the training cost by orders of magnitude. The need for automated tools for improving the quality of supervised learning data is rising as datasets and models are getting larger at an unprecedented speed.
Real-world datasets contain a notable amount of errors Beyer et al. (2020); Northcutt et al. (2021b). Previous works Dau et al. (2022); Nguyen-Duc et al. (2023) showed that removing errors from the training set improves the performance of AI models trained on that dataset. Automatic error rectification, however, is an underexplored topic. In this paper, we present a similarity based approach for error detection and rectification in large scale datasets. We provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Sec. 3.2). Inspired by this observation, we develop simple yet effective similarity-based methods for detecting and rectifying label errors (Sec. 3.3). Extensive experiments demonstrate the superiority of our methods across various settings (Sec. 4.2). Furthermore, our methods are posthoc and model-agnostic i.e., they can be applied to any deep neural network (DNN) architectures without the need for retraining.
2 Background and Related Work
Notation.
[…]
Confident-based approach.
Confident-based methods are based on the notion of confident learning Northcutt et al. (2021a) that deriving label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al. (2022). Low confidence serves as a heuristic indicating the likelihood of label noise. Given a data point $\mathbf{z}$ with output label $\mathbf{y}=(y_{1},...,y_{k},...y_{N})$, the model’s predicted probabilities is $\mathbf{p}=(p_{1},...,p_{k},...,p_{N})$ over $N$ classes. Northcutt et al. (2021a) proposed three label quality scoring methods:
(1) Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$: $\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}}$, for $k\in\{1,2,...,N\}$.
(2) Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class: $\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}}$, for $j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$
(3) Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy: $\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{\mathcal{H}_{N}(\mathbf{p})}$, where $\mathcal{H}_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$.
Gradient-based approach.
Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN. Dau et al. (2022) proposed a way to adapt IF and its variants i.e., Gradient Dot Product (GD; Charpiat et al. (2019a)), Gradient Cosine (GC; Charpiat et al. (2019a)), and Tracing Gradient Decent (TracIn; Pruthi et al. (2020)), for identifying errors in large-scale source code datasets. The idea is that the gradients of error data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points. The algorithm computes the influence score of each data point in the noisy dataset with data points in a reference set. A more negative influence score means it is more likely to be an error. Nguyen-Duc et al. (2023) use class information to improve the performance and stability of these gradient methods.
(1) $\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=-\frac{1}{n}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\mathcal{H}_{\hat{\theta}}^{-1}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})$,
(2) $\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\langle\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})\rangle$,
(3) $\textnormal{GC}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\textnormal{cos}(\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)}))$,
(4) $\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\sum_{t=1}^{T}\eta_{t}\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$, where $\mathcal{H}_{\hat{\theta}}$ is the hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ is the learning rate at epoch $t$.
Other approaches.
The rule-based approach Chu et al. (2013) and statistics-based approach Huang and He (2018) are commonly used for structured data such as tabular data. These methods are not suitable for deep learning, as they assume convexity in the model, and the rules in many large-scale datasets are not easy to find and describe.
3 Method
3.1 Observation
We design experiments to randomly corrupt and inject noise into datasets. We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points. As an illustrative example in Fig. 2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points.
(a) Distribution of (a) Cosine similarity and (b) Dot product over IMDB Maas et al. (2011) with $10\%$ noise. Blue bars represent the similarity between mislabeled data points and their true class data points, red bars represent the similarity between mislabeled data points and other class data points. Features are obtained from a trained BERT model.
We find that this phenomenon persists across varying percentages of noise. To complement our observation, we provide a theoretical analysis in Section 3.2.
3.2 Theoretical Analysis
We consider a deep network receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension. Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias. Given an output function $\phi_{W}$ parameterized by $W\in\mathbb{R}^{N\times d}$. For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we have two output vectors $\phi_{W}(\mathbf{x}^{(i)})$ and $\phi_{W}(\mathbf{x}^{(j)})$ respectively. As seen by a deep network, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{W}(\mathbf{x}^{(i)})$ change would change $\phi_{W}(\mathbf{x}^{(j)})$ as well. If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have large effect on $\phi_{W}(\mathbf{x}^{(j)})$. Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have small effect on $\phi_{W}(\mathbf{x}^{(j)})$. To measure the similarity between data points, we employ a symmetric kernel proposed by Charpiat et al. (2019b): The Inner Product
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{W}\phi_{W}(\mathbf{x}^{(i)}),\nabla_{W}\phi_{W}(\mathbf{x}^{(j)})\right\rangle,$ (1)
We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(W\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$. We have $\phi_{W}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. For simplicity, we remove the bias term. Let denote $\mathbf{u}=W\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. Using the chain rule:
$\nabla_{W}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial W}\right)$
$=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial W}\right)$
$=\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})\mathbf{x}^{\top},$ (2)
[…]
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{\mathbf{y}}}\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}$ (3)
The first term on the right-hand side of Eqn. 3 is the partial derivatives of the loss w.r.t the predicted output $\hat{\mathbf{y}}$. Regarding to the fact: $\mathbf{y}$ is the one-hot vector present label $k$ has element $y_{k}=1$ and $y_{i}=0$ if $i\neq k$. We have:
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{\mathbf{y}}}$ $=\begin{bmatrix}\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{1}}\cdots\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\cdots\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{N}}\end{bmatrix}$
$=\begin{bmatrix}0\cdots\frac{1}{\hat{y}_{k}}\cdots 0\end{bmatrix}$ (4)
The second term is the matrix comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t $\mathbf{u}$. Regarding to the fact:
$\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial u_{1}}&\frac{\partial\hat{y}_{1}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{1}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{k}}{\partial u_{1}}&\frac{\partial\hat{y}_{k}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{k}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{N}}{\partial u_{1}}&\frac{\partial\hat{y}_{N}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{N}}{\partial u_{N}}\\ \end{bmatrix}$ (5)
Substitute Eqn. 5 and Eqn. 4 into Eqn. 3, we get
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=$
$\begin{bmatrix}\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{1}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{k}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{N}}\end{bmatrix}$ (6)
Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{k}$:
$\frac{\partial\hat{y}_{k}}{\partial u_{k}}$ $=\frac{\partial}{\partial u_{k}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$
$=\frac{e^{u_{k}}\left(\sum_{i=1}^{N}e^{u_{i}}\right)-e^{u_{k}}\cdot e^{u_{k}}}{\left(\sum_{i=1}^{N}e^{u_{i}}\right)^{2}}$
$=\hat{y}_{k}(1-\hat{y}_{k})$ (7)
Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{i}$:
$\frac{\partial\hat{y}_{k}}{\partial u_{i}}=\frac{\partial}{\partial u_{i}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$ (8)
Using the chain rule, we get:
$\frac{\partial\hat{y}_{k}}{\partial u_{i}}=-\frac{e^{u_{k}}\cdot e^{u_{i}}}{\left(\sum_{i=1}^{N}e^{u_{i}}\right)^{2}}=-\hat{y}_{k}\hat{y}_{i}$ (9)
Substitute Eqn. 7 and Eqn. 9 into Eqn. 6, we get a column vector:
$\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})=\begin{bmatrix}-\hat{y}_{1}\cdots 1-\hat{y}_{k}\cdots-\hat{y}_{N}\end{bmatrix}^{\top}$ (10)
The inner product kernel in Eqn. 2 become:
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=$ $\overbrace{\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)})^{\top}\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(j)},\mathbf{y}^{(j)})}^{\mathbf{G}^{(ij)}}$
$\cdot(\mathbf{x}^{(j)\top}\mathbf{x}^{(i)})$ (11)
We denote $\mathbf{G}^{(ij)}$ as the dot product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$. Suppose input $\mathbf{x}^{(i)}$ and input $\mathbf{x}^{(j)}$ have label $k$ and $k^{\prime}$ corresponding. We have
$\mathbf{G}^{(ij)}$ $=\begin{bmatrix}-\hat{y}^{(i)}_{1}\cdots 1-\hat{y}^{(i)}_{k}\cdots-\hat{y}^{(i)}_{N}\end{bmatrix}\begin{bmatrix}-\hat{y}^{(j)}_{1}\\ \vdots\\ 1-\hat{y}^{(j)}_{k^{\prime}}\\ \vdots\\ -\hat{y}^{(j)}_{N}\end{bmatrix}$ (12)
we consider 2 cases:
If $k=k^{\prime}$:
$\mathbf{G}^{(ij)}_{k=k^{\prime}}$ $=\hat{y}^{(i)}_{1}\hat{y}^{(j)}_{1}+\cdots+(1-\hat{y}^{(i)}_{k})(1-\hat{y}^{(j)}_{k})$
$+\cdots+\hat{y}^{(i)}_{N}\hat{y}^{(j)}_{N}$
$=(1-\hat{y}^{(i)}_{k})(1-\hat{y}^{(j)}_{k})+\sum_{n=1,n\neq k}^{N}\hat{y}^{(i)}_{n}\hat{y}^{(j)}_{n}$ (13)
If $k\neq k^{\prime}$:
$\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}$ $=\hat{y}^{(i)}_{1}\hat{y}^{(j)}_{1}+\cdots+(1-\hat{y}^{(i)}_{k})(-\hat{y}^{(j)}_{k})$+
$(1-\hat{y}^{(j)}_{k^{\prime}})(-\hat{y}^{(i)}_{k^{\prime}})+\cdots+\hat{y}^{(i)}_{N}\hat{y}^{(j)}_{N}$
$=\hat{y}^{(j)}_{k}(\hat{y}^{(i)}_{k}-1)+\hat{y}^{(i)}_{k^{\prime}}(\hat{y}^{(j)}_{k^{\prime}}-1)$
$+\sum_{n=1,n\neq k,n\neq k^{\prime}}^{N}\hat{y}^{(i)}_{n}\hat{y}^{(j)}_{n}$ (14)
During the training process, the model is more confident about the labels of data points, indicating the value of $\hat{y}^{(i)}_{k}$ and $\hat{y}^{(j)}_{k^{\prime}}$ being closer to $1$. Assume a well-trained model, and $\hat{y}^{(i)}_{k}\approx\hat{y}^{(j)}_{k^{\prime}}=\alpha$; $\hat{y}^{(i)}_{n}=\hat{y}^{(j)}_{n}\approx\epsilon=\frac{1-\alpha}{N-1}$. ($n\neq k$ and $n\neq k^{\prime}$). Substitute these values into Eqn. 13 and Eqn. 14, we get :
$\mathbf{G}^{(ij)}_{k=k^{\prime}}$ $\approx(1-\alpha)^{2}+\epsilon^{2}(N-1)$ (15)
$\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}$ $\approx-\frac{N(1-\alpha)^{2}}{(N-1)^{2}}=-\epsilon^{2}N$ (16)
[…]
$\frac{\mathcal{K}^{(ij)}_{k=k^{\prime}}}{\mathcal{K}^{(ij)}_{k\neq{k^{\prime}}}}=\frac{|\mathbf{G}^{(ij)}_{k=k^{\prime}}|}{|\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}|}$ $\approx\frac{|(1-\alpha)^{2}+\epsilon^{2}(N-1)|}{|-\epsilon^{2}N|}$
$\approx\frac{\epsilon^{2}(N-1)^{2}+\epsilon^{2}(N-1)}{\epsilon^{2}N}$
$\approx N-1$ (17)
[…]
3.3 Algorithm
[…]
Error Detection.
Given a noisy dataset $\mathcal{D}$, for each data point $\mathbf{z}^{(i)}\in\mathcal{D}$ with label $\mathbf{y}^{(i)}$, our algorithm finds $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ such that every data point in $\mathcal{D}$ but not in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ is at most similar to $\mathbf{z}^{(i)}$ as the least similar point in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 8). We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9). The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds. The lower $\mathbf{s}$ is, the more likely a label error. We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11).
Error Rectification.
We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$. Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
Algorithm 1 Error Detection and Rectification
1:
[…]
5:$k$: number of most similar data points.
6:noisy data points in $\mathcal{D}$ are rectified.
7:for $\mathbf{z}^{(i)}\in\mathcal{D}$ do
8:  $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})=\{\mathbf{z}^{(j)}\in\mathcal{D}_{\textnormal{aux}}\}$ s.t. $|\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})|=k$, and
$\sigma(\mathbf{z}^{(i)},\mathbf{z}^{(j)})\geq\max_{\mathbf{z}^{\prime(i)}\in\mathcal{D}\backslash\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})}\sigma(\mathbf{z}^{(i)},\mathbf{z}^{\prime(i)})$
9:  $\mathbf{s}^{(i)}=\frac{1}{k}\sum_{\mathbf{z}^{(j)}\in\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})}\mathbb{I}(\mathbf{y}^{(j)}=\mathbf{y}^{(i)})$
10:end for
[…]
12:for $\mathbf{z}^{(i)}\in\mathcal{D}^{\uparrow}_{:p}$ do
13:  $\mathbf{z}^{(i)}=\left(\mathbf{x}^{(i)},\textsc{Mode}(\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)}))\right)$
14:end for
15:return $\mathcal{D}$
4 Experiment
4.1 Experiment Setting
(a) Error detection accuracy of methods measure on Snippets. The x-axis in the figures presents the change of $t$ from $10\to 100\%$.
Table 1: Test accuracy after removing/rectifying potential noise samples on Snippets with 20% noise. The best and runner-up are marked.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 88.64 (+0.00) 82.50 (+0.00) 79.38 (+0.00)
Confident-W. Entropy 83.02 (-5.62) 84.38 (+1.88) 77.50 (-1.88)
Normalize-Margin 87.19 (-1.45) 87.01 (+4.51) 77.89 (-1.49)
Self-Confidence 86.05 (-2.59) 87.36 (+4.86) 78.02 (-1.36)
Influence Function 83.24 (-5.40) 81.71 (-0.79) 79.51 (+0.13)
Gradient-Cosine 73.81 (-14.83) 83.33 (+0.83) 82.23 (+2.85)
Gradient-Dot 86.53 (-2.11) 82.01 (-0.49) 78.68 (-0.70)
TracIn 85.48 (-3.16) 81.71 (-0.79) 76.67 (-2.71)
Sim-Cos 89.43 (+0.79) 87.85 (-0.79) 85.00 (+2.50) 83.73 (+1.23) 81.53 (+2.15) 83.38 (+4.00)
Sim-Dot 86.71 (-1.93) 87.32 (-1.32) 82.98 (+0.48) 83.95 (+1.45) 81.53 (+2.15) 81.97 (+2.59)
Dataset and model.
We evaluate our method on two common benchmarks: Sentiment Analysis on IMDB Maas et al. (2011) and Topic Classification on Snippets Phan et al. (2008). We use BERT Devlin et al. (2019) as the standard model.
Modeling Realistic Noise.
We construct three realistic, human-originated types of noise: (1) Uniform noise: we randomly select data points and change the label to a different class. (2) Systematic ambiguity noise: we establish a rule $h$, which maps data points in a specific class to another fixed one. This means that the labels of selected instances in class $i$ are flipped to $h(i)$. To ensure distinctiveness, the mapping function $h$ adheres to the condition $h(i)\neq h(j)\forall i,j=\{1,...,N\}$, and $i\neq j$. This noise models the situations where inputs from multiple annotators are often aggregated; the resulting differences in annotations can serve as a model of systematic noise derived from human disagreements. (3) Concentrated noise: we select data points that are densely clustered and change their labels to target labels. We simulate scenarios where the datasets are poisoned by malicious to evaluate the sanitization ability of methods against data poisoning attacks.
Setting.
For each dataset, we construct groups of various sizes of noisy samples by corrupting the label of $p\%$ of the original training data. We construct the auxiliary dataset $\mathcal{D}_{\textnormal{aux}}$ by randomly selecting $m$ samples from the validation set. We fine-tune BERT on the noisy dataset $\mathcal{D}$ and select the best checkpoint measure on the validation set. We select the top $t\%$ ranked samples in $\mathcal{D}^{\uparrow}_{:p}$ and use error detection accuracy for evaluation. After rectifying/removing ranked samples (potentially noisy samples), we re-train the model and report the test accuracy and error reduction rate. Details of the dataset, model, and implementations are in Appendix A.
4.2 Main result
Error detection accuracy.
(1) Fig. 5(a) shows the error detection accuracy of methods with three types of noise with different percentages. As a result, when $t$ increases, the performance of gradient-based methods drastically decreases. This pattern is observed in all three types of noise across different percentages and in both Snippets (Fig. 9) and IMDB (Fig. 10). This result shows that the gradient-based methods are unstable and less inconsistent. (2) Confident-based methods are precise with uniform noises and systematic ambiguity noise, yet struggle with concentrated noise (Fig. 5(a)c). (3) Sim-Cos and Sim-Dot have high detection accuracy and slightly decrease when $t$ increases with different noise. This confirmed that the Similarity-based methods are effective and more robust to ambiguity and concentrated noises than gradient-/confident- based methods. (4) We observe that Sim-Dot often has low detection accuracy on IMDB (Fig. 10). We theoretically proved that for classification datasets with $N$ classes, the similarity of within-class data points is approximately N1N-1 times larger than that of other class data points. For IMDB, where $N=2$, this fraction becomes approximately $1$. That explains why Sim-Dot does not work well on IMDB. Mathematical details are in Section 3.2. (5) Sim-Cos consistently outperforms Sim-Dot by a large margin for all settings. We explain from the standpoint of feature normalization. By definition, Cosine similarity can be seen as the normalized Dot product. In Fig. 9 (b), we empirically show that the noisy samples have $L_{2}$-norm smaller than that of normal samples. Therefore, when dividing the feature of data points by their norm, the similarity between noisy and normal data points tends to be larger, leading to a more distinct distribution of similarities.
Improving datasets’ quality and model generalization.
Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise. Nevertheless, a counterintuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high. We posit that deep models are robust to massive random noise Rolnick et al. (2017), then as the training process continues, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller. The effect of noise, therefore, also decreases as the model converges. When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
5 Conclusion
We introduce similarity-based algorithms for detecting and rectifying errors on large-scale datasets. We theoretically show that the similarity between the penultimate feature’s data points is useful for detecting errors. Experiment results demonstrated the superior performance of our methods, and their capability to improve the dataset quality and model generalization.
Limitations
We discuss the limitations of similarity-based methods: (1) The optimal detection accuracy of Sim-Cos and Sim-Dot unfortunately based on empirical validation, and depends on the choice of $k$ and $\mathcal{D}_{\textnormal{aux}}$ and also with different datasets and model architectures. (2) The generalization of models under the removal or rectification of noise remains uncertain, due to the limited exploration of datasets.
Ethics Statement
We consider only the public datasets and create artificial noises for evaluation. We do not pose any concern about the quality of the original datasets.
Appendix A Implementation detail
A.1 Dataset and model
Snippets Phan et al. (2008)
is an open dataset of web search snippets retrieved from Google Search with 8 domains including Business, Computers, Culture-Arts-Entertainment, Education-Science, Engineering, Health, Politics-Society, and Sports. The training data and testing data include $10,060$ and $2,280$ snippets respectively. For validation purposes, we randomly split original training data into train/valid with the ratio of $8048/2012$. The dataset can be found at http://jwebpro.sourceforge.net/data-web-snippets.tar.gz.
IMDB Maas et al. (2011)
is the most common benchmark for Sentiment Analysis task. IMDB includes $50,000$ reviews from the Internet Movie Database website with original $25,000$ negative and $25,000$ positive reviews. For validation purposes, we randomly split into training, validation, and test sets of sizes $15,000$, $5,000$, and $25,000$. The IMDB dataset can be found at https://ai.stanford.edu/˜amaas/data/sentiment/
BERT Devlin et al. (2019)
stands for Bidirectional Encoder Representations from Transformers. BERT is one of the most standard used pre-trained model for language understanding tasks. In all settings, we use BERT base uncased version.
A.2 Experiment detail
BERT was trained with AdamW Loshchilov and Hutter (2019) with learning rate $\eta=5e-5$, momentum $\beta=(0.9,0.999)$, cross entropy loss, batch-size of $16$ with $15$ epochs. For regularization, we use Dropout Srivastava et al. (2014) of $0.2$. We choose $p=\{5\%,10\%,20\%\}$, $m=1000$, $k=\{1,2,5,10,20,50,100,200\}$, $t=\{10\%,20\%,...,100\%\}$, and $\tau=0.8$. We compute the IF score for BERT with the last layer gradient as previous works Pezeshkpour et al. (2021); Hanawa et al. (2021) and use LiSSA Agarwal et al. (2017) to approximate the Hessian. For TracIn method, we calculate the influence score from the first epoch to the best epoch. We run experiments on $4$ seeds = $\{16,32,64,128\}$ and aggregate the results by taking the mean of these $4$ seeds. Two Nvidia A40s were used to run experiments.
Appendix B Additional results
B.1 Improving datasets and model generalization.
Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise. Nevertheless, a counter-intuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high. We posit that deep models are robust to massive random noise Rolnick et al. (2017), then as the training process, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller. The effect of noise, therefore, also decreases as the model converges. When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
B.2 Effects of Hyperparameters
The effect of the size of $\mathcal{D}_{\textnormal{aux}}$.
We change the size of $\mathcal{D}_{\textnormal{aux}}$ from $100$ to $1500$, fix $k=100$. Tab. 2 and Tab. 3 show the change in error detection accuracy as the size of $\mathcal{D}_{\textnormal{aux}}$ changes. We observed that: (1) the detection accuracy of the methods increases as the size of $\mathcal{D}_{\textnormal{aux}}$ increases. This is true for both Snippets and IMDB, Sim-Cos and Sim-Dot, and for all levels of noise $\{5\%,10\%,20\%\}$. (2) The accuracy gains are larger for smaller $\mathcal{D}_{\textnormal{aux}}$ values. (3) Sim-Cos outperforms Sim-Dot for all noise levels and most $\mathcal{D}_{\textnormal{aux}}$ values.
Table 2: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$, setting with Snippets.
Method 5% noise 10% noise 20% noise
Sim-Cos@100 $16.91$ $22.63$ $31.69$
Sim-Cos@200 $51.99$ $40.67$ $65.25$
Sim-Cos@300 $57.46$ $56.96$ $80.23$
Sim-Cos@400 $60.45$ $72.38$ $86.45$
Sim-Cos@500 $75.37$ $78.60$ $86.88$
Sim-Cos@600 $77.36$ $78.60$ $89.93$
Sim-Cos@700 $82.58$ $78.48$ $93.10$
Sim-Cos@800 $85.07$ $78.60$ $92.91$
Sim-Cos@900 $86.07$ $79.72$ $93.16$
Sim-Cos@1000 $86.56$ $80.01$ $93.28$
Sim-Cos@1500 $88.06$ $80.97$ $94.28$
Sim-Dot@100 $16.91$ $22.63$ $31.69$
Sim-Dot@200 $51.74$ $40.92$ $64.26$
Sim-Dot@300 $57.96$ $56.59$ $79.42$
Sim-Dot@400 $61.19$ $70.64$ $86.82$
Sim-Dot@500 $74.62$ $78.23$ $87.01$
Sim-Dot@600 $76.36$ $78.10$ $88.87$
Sim-Dot@700 $81.34$ $77.36$ $92.91$
Sim-Dot@800 $82.83$ $77.36$ $92.54$
Sim-Dot@900 $83.83$ $78.10$ $92.17$
Sim-Dot@1000 $84.08$ $78.73$ $92.41$
Sim-Dot@1500 $85.32$ $80.34$ $93.78$
Table 3: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$ on detection performance with IMDB.
Method 5% noise 10% noise 20% noise
Sim-Cos@100 $6.40$ $8.55$ $20.10$
Sim-Cos@200 $58.10$ $74.00$ $78.55$
Sim-Cos@300 $60.30$ $75.60$ $78.55$
Sim-Cos@400 $60.10$ $75.65$ $78.55$
Sim-Cos@500 $60.20$ $75.55$ $78.55$
Sim-Cos@600 $59.90$ $75.70$ $78.52$
Sim-Cos@700 $60.20$ $75.55$ $78.45$
Sim-Cos@800 $60.10$ $75.60$ $78.47$
Sim-Cos@900 $60.20$ $75.60$ $78.42$
Sim-Cos@1000 $59.90$ $75.50$ $78.52$
Sim-Cos@1500 $60.10$ $75.55$ $78.50$
Sim-Dot@100 $6.40$ $8.55$ $20.10$
Sim-Dot@200 $57.80$ $66.65$ $77.72$
Sim-Dot@300 $58.00$ $65.70$ $77.80$
Sim-Dot@400 $58.10$ $68.20$ $78.47$
Sim-Dot@500 $57.90$ $67.10$ $77.95$
Sim-Dot@600 $58.00$ $73.65$ $78.35$
Sim-Dot@700 $58.50$ $69.45$ $78.52$
Sim-Dot@800 $57.90$ $65.70$ $77.87$
Sim-Dot@900 $58.20$ $68.65$ $77.47$
Sim-Dot@1000 $59.90$ $68.60$ $77.55$
Sim-Dot@1500 $57.90$ $68.05$ $78.10$
The effect of $k$.
Tab. 4 and Tab. 5 show that Sim-Cos and Sim-Dot have accuracy increase as $k$ increases for all noise levels $\{5\%,10\%,20\%\}$ and for most values of $k$. The accuracy gains are often larger for smaller values of $k$. Sim-Cos outperforms Sim-Dot, especially for larger values of $k$ for all noise levels and most values of $k$. The impact of noise on detection accuracy with Snippets is generally small, but it increases with $k$. For Sim-Cos with $20\%$ noise in Tab. 4, the accuracy starts to decrease after $k=20$. This suggests that using larger $k$ can harm the accuracy when the data is very noisy. For Sim-Dot with $10\%$ and $20\%$ noise in Tab. 4, the accuracy changes are less consistent across different values of $k$. This suggests that Sim-Dot may be more sensitive to the choice of $k$.
Table 4: Error detection accuracy of Sim-Cos and Sim-Dot changes as $k$ changes with Snippets.
Method 5% noise 10% noise 20% noise
Sim-Cos@k=1 $77.61$ $82.46$ $87.69$
Sim-Cos@k=2 $88.55$ $85.57$ $92.23$
Sim-Cos@k=5 $89.30$ $86.31$ $94.53$
Sim-Cos@k=10 $89.30$ $86.94$ $94.65$
Sim-Cos@k=20 $89.55$ $86.69$ $94.53$
Sim-Cos@k=50 $88.06$ $82.09$ $94.46$
Sim-Cos@k=100 $86.56$ $80.10$ $93.28$
Sim-Cos@k=200 $75.12$ $78.23$ $86.76$
Sim-Dot@k=1 $80.09$ $82.09$ $88.75$
Sim-Dot@k=2 $81.34$ $83.45$ $90.05$
Sim-Dot@k=5 $82.58$ $84.70$ $91.29$
Sim-Dot@k=10 $84.57$ $85.32$ $92.60$
Sim-Dot@k=20 $85.57$ $84.57$ $93.10$
Sim-Dot@k=50 $86.07$ $81.96$ $93.59$
Sim-Dot@k=100 $84.08$ $78.73$ $92.41$
Sim-Dot@k=200 $75.12$ $77.73$ $86.76$
Table 5: The effect of $k$ on detection performance with IMDB.
Method 5% noise 10% noise 20% noise
Sim-Cos@k=1 $28.50$ $50.85$ $66.15$
Sim-Cos@k=2 $49.10$ $66.80$ $69.72$
Sim-Cos@k=5 $57.80$ $72.05$ $75.65$
Sim-Cos@k=10 $59.50$ $74.10$ $76.77$
Sim-Cos@k=20 $59.50$ $74.90$ $77.42$
Sim-Cos@k=50 $60.20$ $75.55$ $78.42$
Sim-Cos@k=100 $59.90$ $75.50$ $78.52$
Sim-Cos@k=200 $60.00$ $75.45$ $78.57$
Sim-Dot@k=1 $57.50$ $67.25$ $75.17$
Sim-Dot@k=2 $57.00$ $67.30$ $75.30$
Sim-Dot@k=5 $57.20$ $67.95$ $75.80$
Sim-Dot@k=10 $57.10$ $67.75$ $76.00$
Sim-Dot@k=20 $57.10$ $68.15$ $76.27$
Sim-Dot@k=50 $57.90$ $68.55$ $76.95$
Sim-Dot@k=100 $59.90$ $68.60$ $77.55$
Sim-Dot@k=200 $58.20$ $68.25$ $78.40$
The effect of $\tau$ on error reduction rate.
We analyze the effect of $\tau$ on the error reduction rate. We vary the number of $\tau=\{0.5,0.6,0.7,0.8,0.9,0.99\}$. From results in Tab. 6 and Tab. 7, we see that for both Snippets and IMDB, the change of $\tau$ has minimal impact on the error reduction rate. Generally, the reduction rate is higher when the noise level is higher.
Table 6: Error Reduction Rate on Snippets.
Method 5% noise 10% noise 20% noise
[…]
Table 7: Error Reduction Rate on IMDB.
Method 5% noise 10% noise 20% noise
[…]
B.3 The correlation between detection methods.
We calculate the Spearman correlation between ranking scores assigned to samples by detection methods. Fig. 9(a)a shows that gradient-based, confident-based, and similarity-based methods have low Spearman correlation with each other. Confident-based and Gradient-based methods have negative Spearman correlation, indicating they are very different in ranking. We observed the same phenomenon in datasets across levels and types of noise.
(a) Visualization of (a) Spearman correlation, (b) the $L_{2}$-norm of noisy data points (presented by blue bars) and normal data points (presented by red bars). Using normalization (Cosine; (d)) enables the differentiation of the similarity distribution, therefore, enhancing detection accuracy compared to the unnormalized approach (Dot; (c)).
Table 8: Test accuracy on Snippets with 10% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 85.43 (+0.00) 88.24 (+0.00) 85.08 (+0.00)
Confidence-W. Entropy 88.07 (+2.64) 85.04 (-3.20) 84.56 (-0.52)
Normalize-Margin 81.75 (-3.68) 86.09 (-2.15) 79.73 (-5.35)
Self-Confidence 84.47 (-0.96) 87.06 (-1.18) 82.06 (-3.02)
Influence Function 90.04 (+4.61) 85.92 (-2.32) 81.36 (-3.72)
Gradient-Cosine 82.94 (-2.49) 81.18 (-7.06) 80.48 (-4.60)
Gradient-Dot 86.01 (+0.58) 85.13 (-3.11) 83.29 (-1.79)
TracIn 85.74 (+0.31) 86.88 (-1.36) 77.98 (-7.10)
Sim-Cos 79.73 (-5.70) 87.85 (+2.42) 85.48 (-2.76) 85.57 (-2.67) 81.58 (-3.23) 82.06 (-3.02)
Sim-Dot 88.29 (+2.86) 87.32 (+1.89) 84.34 (-3.90) 86.58 (-1.66) 81.62 (-3.46) 80.83 (-4.25)
Table 9: Test accuracy on Snippets with 5% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 87.14 (+0.00) 87.36 (+0.00) 87.93 (+0.00)
Confidence-W. Entropy 84.86 (-2.28) 86.57 (-0.79) 87.80 (-0.13)
Normalize-Margin 84.65 (-2.49) 84.82 (-2.54) 83.77 (-4.16)
Self-Confidence 85.35 (-1.79) 83.99 (-3.37) 86.09 (-1.84)
Influence Function 85.30 (-1.84) 86.00 (-1.36) 85.30 (-2.63)
Gradient-Cosine 85.43 (-1.71) 85.13 (-2.23) 79.51 (-8.42)
Gradient-Dot 84.56 (-2.58) 89.21 (+1.85) 87.06 (-0.87)
TracIn 84.65 (-2.49) 87.71 (+0.35) 81.88 (-6.05)
Sim-Cos 86.45 (-0.69) 85.17 (-1.97) 87.10 (-0.26) 85.48 (-1.88) 85.39 (-2.54) 87.19 (-0.74)
Sim-Dot 87.15 (+0.01) 87.06 (-0.08) 83.37 (-3.99) 84.34 (-3.02) 84.12 (-3.81) 79.82 (-8.11)
Table 10: Test accuracy on IMDB with 20% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 89.92 89.36 85.98
Confidence-W. Entropy 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Normalize-Margin 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Self-Confidence 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Influence Function 90.95 (+1.03) 90.51 (+1.15) 87.09 (+1.11)
Gradient-Cosine 90.59 (+0.67) 90.93 (+1.57) 86.78 (+0.80)
Gradient-Dot 90.43 (+0.51) 89.45 (+0.09) 82.87 (-3.20)
TracIn 90.19 (+0.27) 91.09 (+1.73) 82.87 (-3.20)
Sim-Cos 90.70 (+0.78) 88.73 (-1.19) 90.78 (+1.42) 90.83 (+1.47) 87.76 (+1.78) 84.73 (-1.25)
Sim-Dot 91.49 (+1.57) 90.46 (+0.54) 91.58 (+2.22) 90.13 (+0.77) 87.25 (+1.27) 87.46 (+1.48)
Table 11: Test accuracy on IMDB with 10% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 91.76 (+0.00) 91.68 (+0.00) 89.41 (+0.00)
Confidence-W. Entropy 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Normalize-Margin 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Self-Confidence 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Influence Function 91.65 (-0.11) 90.97 (-0.71) 86.26 (-3.15)
Gradient-Cosine 91.72 (-0.04) 91.29 (-0.39) 88.05 (-1.36)
Gradient-Dot 90.81 (-0.95) 92.10 (+0.42) 88.15 (-1.26)
TracIn 91.75 (-0.01) 91.66 (-0.02) 88.86 (-0.55)
Sim-Cos 91.12 (-0.64) 91.98 (+0.22) 92.52 (+0.84) 92.17 (+0.49) 90.15 (+0.74) 89.20 (-0.21)
Sim-Dot 91.38 (-0.38) 91.51 (-0.25) 91.52 (-0.16) 91.61 (-0.07) 89.78 (+0.37) 88.37 (-1.04)
Table 12: Test accuracy on IMDB with 5% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 91.67 (+0.00) 92.07 (+0.00) 91.52 (+0.00)
Confidence-W. Entropy 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Normalize-Margin 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Self-Confidence 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Influence Function 92.55 (+0.88) 91.76 (-0.31) 90.77 (-0.75)
Gradient-Cosine 92.46 (+0.79) 90.88 (-1.19) 89.78 (-1.74)
Gradient-Dot 91.79 (+0.12) 92.58 (+0.51) 90.32 (-1.20)
TracIn 92.62 (+0.95) 91.97 (-0.10) 91.45 (-0.07)
Sim-Cos 92.10 (+0.43) 92.10 (+0.43) 93.09 (+1.02) 92.50 (+0.43) 92.11 (+0.59) 91.74 (+0.22)
Sim-Dot 92.69 (+1.02) 92.06 (+0.39) 91.63 (-0.44) 91.09 (-0.98) 91.59 (+0.07) 90.96 (-0.56)
Figure 9: Detection accuracy of methods measured on Snippets with different levels of noise.
Figure 10: Detection accuracy of methods measured on IMDB with different levels of noise.